The perimeter of a rectangle is 56cm. The width is 8cm. less than the length. Find its area.
Found 2 solutions by math_helper, rumessa:
Answer by math_helper(2461)


P = 2L + 2W
56 = 2L + 2(L-8) <<<— Width is Length-8cm
56 = 2L + 2L - 16
Can you solve it from here? You will need to find L, then W (=L-8) and then compute L*W to get area.

Answer by rumessa(13)

length =l
width = l-8
p = 56
p= 2(l+w)
56 = 2 ( l +l -8 )
56 = 2( 2l -8 )
56 = 4i -16
4l = 56=16
4l = 72
l= 18
w = l - 8
w = 18 - 8
w = 10